OpenStudy (anonymous):

Which statement correctly compares 8/15 and 5/8? A. 8/15>5/8 B. 8/15<5/8 C. 8/15=5/8

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

anyone please help

OpenStudy (jango_in_dtown):

8*8=64 and 15*5=75 75>64 so 15*5>8*8 implies 5/8>8/15

OpenStudy (anonymous):

LCM is 120 2^3 * 3 * 5 8/15 = 64/120 5/8 = 75/120 Since 75/120 > 64/120, --> 5/8 > 8/15

OpenStudy (jango_in_dtown):

i.e. 8/15<5/8